4. A scuba diver with his gear has a mass of 110 kg. When testing his gear in the pool, he must add 9.00 kg of lead to achieve neutral buoyancy (i.e. the forces acting on him are balanced.) How much additional lead must he wear to achieve neutral buoyancy in the ocean? (The density of lead is 11.3 g/cc. Assume the density of pool water is 1.00 g/cc, the density of ocean water is 1.03 g/cc.)
Answer #1

Solution :- Given data,

Mass 110 kg

Density of lead = 11.3 g/cc

Density of pool water = 1.00 g/cc

Density of ocean water = 1.03 g/cc

Buoyant force =volume of fluid displaced*g*density

In pool

(110+9)×9.8=V×1000×9.8

We got V=0.119

In ocean now

Let lead added is x kg

Buoyant force=0.119×9.8×1030

110+x=0.119×1030

Lead required

x=12.57 kg